### v1

# Unleashing the Power of Neural Collapse: Consistent Supervised-Unsupervised Alignment for Generalized Category Discovery

Abstract
Generalized Category Discovery (GCD) focuses on classifying known categories while simultaneously discovering novel categories from unlabeled data. However, previous GCD methods face challenges due to inconsistent optimization objectives and category confusion. This leads to feature overlap and ultimately hinders performance on novel categories. To address these issues, we propose the Neural Collapse-inspired Generalized Category Discovery (NC-GCD) framework. By pre-assigning and fixing Equiangular Tight Frame (ETF) prototypes, our method ensures an optimal geometric structure and a consistent optimization objective for both known and novel categories. We introduce a Consistent ETF Alignment Loss that unifies supervised and unsupervised ETF alignment and enhances category separability. Additionally, a Semantic Consistency Matcher (SCM) is designed to maintain stable and consistent label assignments across clustering iterations. Our method achieves strong performance on multiple GCD benchmarks, significantly enhancing novel category accuracy and demonstrating its effectiveness.
(a) The difference between our method and previous studies.
(b) Illustration of the proposed NC-GCD.
Figure 1: Illustration of the differences between prior studies, our proposed NC-GCD framework, and the overview of the NC-GCD. Our method pre-assigns fixed ETF prototypes rather than dynamically learning prototypes, ensuring consistent optimization objectives for known and novel categories.
1 Introduction
Generalized Category Discovery (GCD) has raised emerging attention in recent years, which aims to simultaneously classify known classes and discover novel categories from unlabeled data. Unlike traditional semi-supervised learning, which assumes all classes are predefined during training, GCD operates in a more realistic open-world scenario [1, 2, dong2024brain,liu2020openvocabulary] where only a small portion of known-class data is labeled, and all novel-class samples lack both labels and category information. It requires models to utilize limited supervision from labeled known classes while autonomously discovering semantically coherent clusters in the unlabeled subset without assuming predefined class structures or a clear distinction between known and novel distributions.
Recent research in GCD has witnessed substantial progress. Early research [3] introduces a contrastive learning framework to distinguish known and novel categories. On this basis, contrastive learning-based methods [4, 5] refine category boundaries using dynamic contrastive learning and prompt-based affinity learning. Additionally, representation learning-based methods[6, 7, 8] focus on optimizing feature representations to enhance category separability.
However, most existing methods dynamically optimize cluster prototypes or classification weights, leading to two key issues. First, Inconsistent Optimization Objective: Existing frameworks lack a consistent optimization objective for known and novel categories. As a result, the models tend to prioritize learning from labeled known categories while failing to establish proper decision boundaries for the novel ones. Second, Category Confusion: As shown in Fig. 1(a), existing methods lack geometric constraints on feature distributions. Moreover, novel categories are optimized entirely in an unsupervised manner, making it even more difficult to achieve sufficient separation between feature-similar categories, leading to category overlap and reduced accuracy. Based on the above observations, we wonder:
Can we pre-assign an optimal geometric structure where both known and novel categories are equally well-separated, enabling consistent learning for all classes by aligning features to this structure?
Neural Collapse (NC) is a phenomenon in well-trained classification networks, where features of each category align with a Simplex Equiangular Tight Frame (ETF) structure [9]. In this structure, within-category features collapse to their respective category mean, and the means of different categories are positioned at the vertices of a simplex. This alignment maximizes inter-category separation while maintaining within-class compactness, creating an optimal geometric arrangement for classification tasks.
Motivated by Neural Collapse theory, we propose the Neural Collapse-inspired Generalized Category Discovery (NC-GCD) framework, which leverages NC principles to create a structured feature space for both known and novel categories. Unlike existing methods that dynamically learn prototypes, as shown in Fig. 1(b), our approach pre-assigns ETF prototypes before training, ensuring an optimal geometric structure and a consistent optimization objective. Our framework integrates three key components: 1) Unsupervised ETF Alignment periodically clustering all categories and aligning the top $\alpha$most confident samples to their respective ETF prototypes. This process enhances feature separability, particularly for novel categories. 2) Supervised ETF Alignment stabilizes known categories by aligning labeled features with their corresponding ETF prototypes. Afterwards, we unify supervised and unsupervised alignment by designing a Consistent ETF alignment loss. However, clustering instability may lead to inconsistent label assignments, where samples of the same category are mapped to different clusters across multiple iterations, thereby disrupting ETF alignment. Even more critically, the direct mapping between ground-truth labels and ETF prototypes in supervised alignment can introduce mismatches, further compromising model stability. 3) To address these issues, we introduce the Semantic Consistency Matcher (SCM), which enforces one-to-one alignment between clusters across iterations, stabilizing pseudo-label assignments and ensuring proper mapping of supervised labels to ETF prototypes. By integrating these components, our method establishes a consistent optimization objective, building a structured feature space that enhances class separability and effectively mitigates category confusion.
Our contributions are summarized as follows:
  • We propose a fixed ETF prototype framework that establishes a consistent optimization objective for both known and novel categories, improving category separability through a consistent supervised-unsupervised alignment.
  • We introduce the Semantic Consistency Matcher to maintain consistency in cluster labels across multiple iterations, thereby stabilizing the training process and reducing clustering-induced fluctuations.
  • Our method, as demonstrated in Fig. 2, achieves state-of-the-art performance on six GCD benchmarks, with significant improvements in novel category accuracy.
Figure 2: Performance Overview. Compared to previous studies, our method exhibits superior overall accuracy, with a notable improvement in the accuracy of novel categories.
2 Related Work
2.1 Generalized Category Discovery (GCD)
Generalized Category Discovery (GCD) identifies known and novel categories within a partially labeled dataset. Early methods[3] leveraged contrastive learning to distinguish categories. Later, DCCL [4] and PromptCAL [5] refined category boundaries using dynamic contrastive learning and prompt-based affinity mechanisms. GPC [10] and SimGCD [11] introduced Gaussian Mixture Models and parametric classification frameworks to estimate category distributions. Other methods have focused on self-supervised clustering and feature refinement. PIM [6], AGCD [8], and CMS [7] introduced contrastive mean-shift learning and representation alignment strategies to improve novel category separability. Meanwhile, approaches like UNO [12] and ORCA [13] explored semi-supervised representation learning, while RankStats [14] employed ranking-based statistics to facilitate unknown category discovery. However, most existing methods rely on dynamically learned prototypes and lack a consistent optimization target. This results in category confusion and imbalanced learning.
2.2 Neural Collapse and its Applications
Neural Collapse (NC) describes a geometric phenomenon in deep networks where, in the final stage of training, category feature representations converge to form a Simplex Equiangular Tight Frame (ETF) structure [9]. This structure ensures optimal category separation and minimal intra-category variance, making it highly effective for classification. Several works have provided theoretical insights into Neural Collapse under different loss functions, demonstrating that it emerges naturally in both cross-entropy loss [15, 16, 17, 18] and mean squared error (MSE) loss [19, 20, 21, 22].
Beyond theoretical exploration, NC has been applied to various machine-learning tasks. In imbalanced learning, NC principles have been leveraged to stabilize classifier prototypes under category-imbalanced training [23, 24]. In few-shot class-incremental learning, studies have explored ETF-based classifiers to enhance feature-classifier alignment and mitigate forgetting [25]. In federated learning, NC has been shown to improve model generalization under non-iid settings [26]. Furthermore, NC-inspired approaches have been developed to address large-scale classification problems, such as optimizing one-vs-rest margins for generalized NC [27]. Our NC-GCD framework extends NC principles by pre-assigning ETF prototypes to known and novel categories. By integrating supervised and unsupervised ETF alignment, our method effectively addresses GCD’s misalignment and category confusion.
3 Preliminaries
3.1 Problem Setting of GCD
Generalized Category Discovery (GCD) aims to identify novel categories in an unlabeled dataset while maintaining classification performance for known categories. Formally, we define a dataset $\mathcal{D}=\mathcal{D}^{l}\cup\mathcal{D}^{u}=\{(x_{i},y_{i})\}$, where the labeled dataset is given by $\mathcal{D}^{l}=\{(x_{i}^{l},y_{i}^{l})\}\subset\mathcal{X}\times\mathcal{Y}^{l}$, containing samples $x_{i}^{l}$ with corresponding labels $y_{i}^{l}$ from the set of known categories $\mathcal{Y}^{l}$. The unlabeled dataset is defined as $\mathcal{D}^{u}=\{x_{i}^{u}\}\subset\mathcal{X}$, which consists of samples from both known and novel categories. The overall category set is $\mathcal{Y}^{u}=\mathcal{Y}^{l}\cup\mathcal{Y}^{n}$, where $\mathcal{Y}^{l}\cap\mathcal{Y}^{n}=\emptyset$ and $\mathcal{Y}^{n}$ represents the unknown novel categories. The goal of GCD is to train a model that can accurately classify samples from $\mathcal{Y}^{l}$ while discovering and organizing samples from $\mathcal{Y}^{n}$.
3.2 Neural Collapse and Definition of Simplex ETF
Neural Collapse, observed during the terminal training phase in well-regularized neural networks on balanced datasets [9], describes a symmetric geometric structure in the last-layer features and classifier weights. Features of the same category collapse to their within-category mean, which aligns with the corresponding classifier weights, forming a Simplex Equiangular Tight Frame (ETF).
Definition of Simplex ETF. A Simplex ETF consists of $K$ vectors in $\mathbb{R}^{d}$ that form an optimal geometric configuration for classification. These vectors, denoted as $P=[p_{1},p_{2},\dots,p_{K}]\in\mathbb{R}^{d\times K}$, satisfy:
$P=\sqrt{\frac{K}{K-1}}U\left(I_{K}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right),$ (1)
where $U\in\mathbb{R}^{d\times K}$ is an orthogonal matrix satisfying $U^{T}U=I_{K}$, $I_{K}$ is the $K\times K$ identity matrix, and $\mathbf{1}_{K}$ is a $K$-dimensional all-ones vector. Each prototype $e_{k}$ has unit $\ell_{2}$-norm ($\|e_{k}\|=1$) and fixed pairwise cosine similarity:
$p_{k}^{\top}p_{j}=\frac{K}{K-1}\delta_{k,j}-\frac{1}{K-1},\quad\forall k,j\in[1,K],$ (2)
where $\delta_{k,j}$ is the Kronecker delta.
Neural Collapse Phenomenon. The neural collapse phenomenon can be summarized as follows:
(NC1) Feature Collapse: Last-layer features of the same category collapse to their within-class mean: $\{\Sigma_{W}^{(k)}\to 0$, $\Sigma_{W}^{(k)}=\text{Avg}\{(\mu_{k,i}-\mu_{k})(\mu_{k,i}-\mu_{k})^{T}\}$ where $\mu_{k,i}$ is the feature of the $i$-th sample in the class $k$ and $\mu_{k}$ is the within-class mean.
(NC2) Convergence to Simplex ETF: The within-class means of all classes, centered by the global mean $\mu_{G}=\text{Avg}_{i,k}(\mu_{k,i})$, converge to the vertices of a Simplex ETF: $\hat{\mu}_{k}={\mu_{k}-\mu_{G}}/{\|\mu_{k}-\mu_{G}\|}$, $\forall k\in[1,K]$.
(NC3) Self-Duality: Within-class means align with the corresponding classifier weights: $\hat{\mu}_{k}={w_{k}}/{\|w_{k}\|}$, $w_{k}$ is the weight of the class $k$.
(NC4) Simplified Prediction: Model prediction simplifies to a nearest-class-center rule: $\hat{y}=\arg\min_{k}\|\mu-\mu_{k}\|=\arg\max_{k}\langle\mu,w_{k}\rangle,$ where $\mu$ is the feature of a test sample.
Figure 3: Overview of the NC-GCD Framework. The framework uses periodic clustering to group features, aligning them with pre-assigned ETF prototypes. The Semantic Consistency Matcher (SCM) ensures consistent label assignments across clustering iterations. This process stabilizes feature alignment and maintains a consistent geometric structure for both known and novel categories.
4 Method
4.1 Overview
As shown in Fig.3, our NC-GCD framework comprises four main components: a pre-trained visual encoder $f(\cdot)$, a periodic clustering module $g(\cdot)$, a pre-assigned ETF prototype set $P$, and the Semantic Consistency Matcher $\phi_{\text{SCM}}(\cdot)$.
First, we utilize the pre-trained visual encoder $f(\cdot)$ to extract the image embedding $e$ and augmented image embedding $e^{\prime}$ for each sample. These embeddings are passed to the periodic clustering module $g(\cdot)$, which groups samples into clusters and computes the clustering prototypes $\{c_{1},c_{2},\dots,c_{K}\}$. On the one hand, unsupervised ETF alignment pulls the top $\alpha$of samples most similar to their cluster centers toward their corresponding pre-assigned ETF prototypes using a dot-regression loss. On the other hand, supervised ETF alignment ensures that labeled samples are aligned with their corresponding ETF prototypes, maintaining geometric consistency.
However, two key challenges arise. First, clustering is inherently unstable, leading to inconsistencies where the same category may be assigned to different clusters across iterations, making it difficult to maintain alignment with pre-assigned ETF prototypes. Second, in supervised alignment, the direct mapping between true labels and ETF prototypes may not always hold, creating mismatches that affect model stability. To address these issues, we introduce the Semantic Consistency Matcher (SCM), which guarantees one-to-one matching between clusters over time, reducing inconsistencies caused by clustering dynamics. By integrating these components, our method effectively mitigates category confusion and label imbalance, ensuring a structured feature space and improved performance in generalized category discovery.
4.2 ETF Alignment
Pre-assigned ETF Prototype. We construct a Simplex Equiangular Tight Frame (ETF) as pre-assigned classifier prototypes to enforce a structured feature space and maximize category separability. Given or estimated a number of categories $K$, we can get its categories set $\mathcal{Y}_{t}$. Then we define the optimal ETF prototypes $\mathbf{P}=\{p_{1},p_{2},\dots,p_{K}\}$ as Eq.1.
Each prototype $p_{i}\in\mathbb{R}^{d}$ maintains unit $\ell_{2}$-norm and follows a fixed pairwise similarity:
$p_{i}^{\top}p_{j}=\frac{K}{K-1}q_{i}^{\top}q_{j}-\frac{1}{K-1},\quad\forall i,j\in[1,K].$ (3)
By pre-assigning the ETF prototypes, we provide a stable and optimal alignment structure for both known and novel categories.
Unsupervised ETF Alignment. For each sample $x_{i}\in\mathcal{D}$, we first extract its image embedding $e_{i}=f(x_{i})$ and augmented embedding $e_{i}^{\prime}$. To introduce structural constraints and stabilize feature learning, we perform periodic clustering on the unlabeled dataset $\mathcal{D}$ every $T$ epochs. Given the extracted image embeddings $\{e_{1},\dots,e_{N}\}$, we apply clustering to obtain $K$ cluster centers $\{c_{1},\dots,c_{K}\}$. Then, each sample embedding $e_{i}$ is assigned to the nearest cluster based on the cosine similarity:
$\hat{y}_{i}=\arg\max_{k}\frac{e_{i}^{\top}c_{k}}{\|e_{k}\|\|c_{k}\|},\quad\forall x_{i}\in\mathcal{D}.$ (4)
Once pseudo-labels $\hat{y}_{i}$ are assigned, we compute the similarity between each sample and its respective cluster center:
$s_{i}=\frac{e_{i}^{\top}c_{k}}{\|e_{i}\|\|c_{k}\|},\quad\forall x_{i}\in\mathcal{D}_{k}.$ (5)
where $\mathcal{D}_{k}$ is the set of samples assigned to cluster k.
To align features with the pre-assigned ETF prototype $\mathbf{P}=\{p_{1},p_{2},\dots,p_{K}\}$, we select the top $\alpha\%$ of samples with the highest similarity $s_{i}$ within each cluster. These high-confidence samples are pulled toward the corresponding ETF prototype $p_{k}$ using a Dot-Regression Loss[23]:
$\mathcal{L}_{\text{ETF}}^{u}=\frac{1}{|\tilde{D}_{k}|}\sum_{e_{i}\in\tilde{D}_{k}}\|e_{i}-p_{k}\|^{2},$ (6)
where $\tilde{D}_{k}=\{x_{i}\mid s_{i}\text{ is in the top }\alpha\%\text{ of }\mathcal{D}_{k}\}$ represents the selected high-confidence samples assigned to cluster $c_{k}$.
Supervised ETF Alignment. For each sample $x_{i}^{l}\in\mathcal{D}^{l}$, we extract its feature embedding $e_{i}^{l}$. The SMC establishes a one-to-one mapping between predicted clusters and ground-truth labels, transforming the original labels $y_{i}^{l}$ into ETF-aligned labels $y_{i}^{\text{{ETF}}}=\phi_{\text{{SCM}}}({y}_{i}^{l})$ (detailed in Section 4.3). Subsequently, each labeled sample is encouraged to align with its assigned ETF prototype:
$\mathcal{L}_{\text{ETF}}^{s}=\frac{1}{|\mathcal{D}^{l}|}\sum_{x_{i}^{l}\in\mathcal{D}^{l}}\|e_{i}^{l}-p_{a}\|^{2},a\in[1,K]$ (7)
where $a=y_{i}^{\text{{ETF}}}$. This loss enforces geometric consistency between known category features and their corresponding ETF prototypes, ensuring a structured feature space.
Consistent ETF Alignment Loss. To balance the contributions of supervised and unsupervised ETF alignment, we define the hybrid ETF loss as:
$\mathcal{L}_{\text{ETF}}=(1-\gamma)\mathcal{L}_{\text{ETF}}^{u}+\gamma\mathcal{L}_{\text{ETF}}^{s},$ (8)
where $\gamma$ is a weighting coefficient determining the trade-off between unsupervised and supervised ETF alignment.
4.3 Semantic Consistency Matcher (SCM)
In periodically clustering-based learning, two key challenges arise: 1) Pseudo-label instability across iterations. Clustering fluctuations introduce inconsistencies, disrupting the learning process and degrading feature alignment. 2) Misalignment between the predicted clusters and the true labels in supervised alignment. Supervised ETF alignment may misassign cluster labels, leading to further optimization instability. To mitigate these issues, we propose SCM, which stabilizes pseudo-label assignments across iterations and ensures consistent alignment between predicted clusters and true labels. By enforcing a one-to-one mapping between clustering results from consecutive iterations, SCM reduces clustering-induced fluctuations.
Specifically, SCM aligns pseudo-labels between consecutive clustering iterations through an optimal assignment strategy. Given prediction sets $\{\hat{y}_{1}^{t},\dots,\hat{y}_{N}^{t}\},\quad\forall\hat{y}_{i}^{t}\in\mathcal{Y}_{t}$ for the current iteration and $\{\hat{y}_{1}^{t-1},\dots,\hat{y}_{N}^{t-1}\},\quad\forall\hat{y}_{i}^{t-1}\in\mathcal{Y}_{t-1}$ for the previous iteration, where $\mathcal{Y}_{t}$ and $\mathcal{Y}_{t-1}$ represent the corresponding label sets. Then we seek an optimal permutation $\sigma^{*}$ that maximizes label consistency:
$\sigma^{*}=\arg\max_{\sigma\in S_{K}}\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{I}(\hat{y}_{i}^{t}=k)\mathbb{I}(\hat{y}_{i}^{t-1}=\sigma(k)),$ (9)
where $\sigma$ is a bi-jective mapping function ensuring a one-to-one correspondence between clusters from consecutive iterations, and $S_{K}$ is the space of all valid permutations over $K$ clusters. The indicator function $\mathbb{I}(\cdot)$ returns 1 if the condition holds and 0 otherwise. The label set in the current iteration is updated as $\mathcal{Y}_{t}=\sigma^{*}(\mathcal{Y}_{t-1})$.
SCM also ensures consistency in supervised ETF alignment by optimally mapping predicted clusters to their corresponding ETF-aligned labels. Given the prediction of the labeled samples and the set of ground truth labels, SCM applies the same optimal assignment strategy to obtain the permutation $\sigma^{l}$. Using this mapping, the ETF-aligned label set is updated as: $\mathcal{Y}_{t}^{\text{{ETF}}}=\sigma^{l}(\mathcal{Y}^{l})$.
4.4 Comparative Representation Learning
The unsupervised loss encourages similarity between embeddings of the same sample under different augmentations while distinguishing them from other samples [7]:
$\mathcal{L}_{\text{rep}}^{u}=-\frac{1}{|B|}\sum_{i\in B}\log\frac{\exp(e_{i}\cdot e_{i}^{\prime}/\tau)}{\sum_{j\neq i}\exp(e_{i}\cdot e_{j}/\tau)},$ (10)
where $e_{i}^{\prime}$ is the augmented view of $e_{i}$, $\tau$ is the temperature parameter, and $|B|$ is the batch size.
For labeled samples, we apply a supervised loss to enforce compact intra-category representations:
$\mathcal{L}_{\text{rep}}^{s}=-\frac{1}{|B_{l}|}\sum_{i\in B_{l}}\frac{1}{|H(i)|}\sum_{h\in H(i)}\log\frac{\exp(e_{i}^{l}\cdot e_{h}/\tau)}{\sum_{j\neq i}\exp(e_{i}^{l}\cdot e_{j}/\tau)},$ (11)
where $H(i)$ represents the set of samples sharing the same label as $i$. The total representation learning objective is:
$\mathcal{L}_{\text{REP}}=(1-\lambda)\mathcal{L}_{\text{REP}}^{u}+\lambda\mathcal{L}_{\text{REP}}^{s}.$ (12)
where $\lambda$ balances the contributions of unsupervised and supervised components.
4.5 Final Objective
The final loss function integrates all components to ensure stable feature alignment and category separability:
$\mathcal{L}=\beta\mathcal{L}_{\text{ETF}}+\mathcal{L}_{\text{REP}},$ (13)
where $\beta$ controls the relative contribution of ETF alignment loss and representation learning loss.
Method CUB Stanford Cars FGVC Aircraft Herbarium 19 CIFAR100 ImageNet100
All Old New All Old New All Old New All Old New All Old New All Old New
(a) Clustering with the ground-truth number of classes $K$ given
GCD (CVPR 2022) 51.3 56.6 48.7 39.0 57.6 29.9 45.0 41.1 46.9 35.4 51.0 27.0 73.0 76.2 66.5 74.1 89.8 66.3
DCCL (CVPR 2023) 63.5 60.8 64.9 43.1 55.7 36.2 - - - - - - 75.3 76.8 70.2 80.5 90.5 76.2
PromptCAL (CVPR 2023) 62.9 64.4 62.1 50.2 70.1 40.6 52.2 52.2 52.3 37.0 52.0 28.9 81.2 84.2 75.3 83.1 92.7 78.3
GPC (ICCV 2023) 55.4 58.2 53.1 42.8 59.2 32.8 46.3 42.5 47.9 - - - 77.9 85.0 63.0 76.9 94.3 71.0
SimGCD (ICCV 2023) 60.3 65.6 57.7 53.8 71.9 45.0 54.2 59.1 51.8 44.0 58.0 36.4 80.1 81.2 77.8 83.0 93.1 77.9
PIM (CVPR 2023) 62.7 75.7 56.2 43.1 66.9 31.6 - - - 42.3 56.1 34.8 78.3 84.2 66.5 83.1 95.3 77.0
SelEx (ECCV2024) 73.6 75.3 72.8 58.5 75.6 50.3 57.1 64.7 53.3 - - - 82.3 85.3 76.3 83.1 93.6 77.8
CMS (CVPR 2024) 68.2 76.5 64.0 56.9 76.1 47.6 56.0 63.4 52.3 36.4 54.9 26.4 82.3 85.7 75.5 84.7 95.6 79.2
SPT (ICLR 2024) 65.8 68.8 65.1 59.0 79.2 49.3 59.3 61.8 58.1 43.4 58.7 35.2 81.3 84.3 75.6 85.4 93.2 81.4
Ours (NC-GCD) 74.8 76.8 73.8 59.9 77.8 51.2 60.0 57.6 61.2 46.4 58.4 40.7 82.7 85.5 77.3 88.4 94.1 85.5
(b) Clustering without the ground-truth number of classes $K$ given
GCD (CVPR 2022) 51.1 56.4 48.4 39.1 58.6 29.7 - - - 37.2 51.7 29.4 70.8 77.6 57.0 77.9 91.1 71.3
GPC (ICCV 2023) 52.0 55.5 47.5 38.2 58.9 27.4 43.3 40.7 44.8 36.5 51.7 27.9 75.4 84.6 60.1 75.3 93.4 66.7
PIM (CVPR 2023) 62.0 75.7 55.1 42.4 65.3 31.3 - - - 42.0 55.5 34.7 75.6 81.6 63.6 83.0 95.3 76.9
CMS (CVPR 2024) 64.4 68.2 62.2 51.7 68.9 43.4 55.2 60.6 52.4 37.4 56.5 27.1 79.6 83.2 72.3 81.3 95.6 74.2
Ours (NC-GCD) 70.3 72.1 69.4 54.0 73.1 44.8 55.4 57.3 54.5 42.3 56.2 34.8 80.5 83.7 74.0 85.7 95.9 80.6
Table 1: Comparison with the state of the arts on GCD, evaluated with or without the GT K for clustering, with DINOv1 backbone.
5 Experiments
5.1 Experimental Setup
Datasets. Our evaluation spans six image classification benchmarks, including both generic and fine-grained datasets. For generic datasets, we use CIFAR-100 [28] and ImageNet-100 [29, 30]. For fine-grained datasets, we evaluate on CUB-200 [31], Stanford Cars [32], FGVC Aircraft [33], and Herbarium19 [34]. To separate classes into known and novel categories, we follow the SSB split protocol [35] for the fine-grained datasets. For CIFAR-100 and ImageNet-100, we perform a random class split using a fixed seed, consistent with previous studies [36]. More details regarding the class splits can be found in the Appendix.
Implementation Details. Our method utilizes the pre-trained DINO ViT-B/16 [37, 38] as the backbone network. We fine-tune the final layer of the image encoder along with a projection head, following prior work [3, 7]. The projection head consists of an MLP with a 768-dimensional input, a 2048-dimensional hidden layer, and a 768-dimensional output, followed by a GeLU activation function [39]. The learning rate is set to 0.1. Other hyperparameters, including batch size, temperature $\tau_{s}$, weight decay $\lambda$, and the number of augmentations, are set to 128, 0.07, $1e^{-4}$, 0.35, and 2, respectively, in accordance with previous studies [36]. All experiments are conducted on an NVIDIA A6000 GPU.
Method Fine-grained datasets Avg Classification datasets Avg All datasets Avg
All Old New All Old New All Old New
(a) Clustering with the ground-truth number of classes K given
SimGCD (ICCV 2023) 53.1 63.7 47.7 81.6 87.2 77.9 62.6 71.5 57.8
CMS (CVPR 2024) 54.4 67.7 47.6 83.5 90.7 77.4 64.1 75.4 57.5
SPT (ICLR 2024) 56.9 67.1 51.9 83.4 88.8 78.5 65.7 74.3 60.8
Ours 60.3 67.6 56.7 85.5 89.8 81.4 68.7 75.0 64.9
Improv. over SPT +3.4 +0.5 +4.8 +2.2 +1.0 +2.9 +3.0 +0.7 +4.2
(b) Clustering without the ground-truth number of classes K given
GPC (ICCV 2023) 46.2 50.8 38.9 75.4 89.0 63.4 53.5 64.1 45.7
CMS (CVPR 2024) 52.2 63.6 46.3 80.5 89.4 73.3 61.6 72.2 55.3
Ours 55.5 64.6 50.9 83.1 89.8 77.3 64.7 73.0 59.7
Improv. over CMS +3.3 +1.1 +4.6 +2.6 +0.4 +4.1 +3.1 +0.9 +4.4
Table 2: Comparison of various methods across multiple metrics, including Fine-grained datasets, Classification datasets, and All datasets.
Sup ETF Alignment Unsup ETF Alignment CUB Stanford Cars FGVC Aircraft Herbarium 19 CIFAR100 ImageNet100 Avg
All Old New All Old New All Old New All Old New All Old New All Old New All Old New
67.7 75.7 63.9 55.1 74.3 45.8 54.3 57.4 52.8 36.5 55.0 26.5 82.0 85.4 75.2 84.4 94.0 80.8 63.3 73.7 57.5
69.6 75.8 66.5 57.0 75.1 46.7 56.1 62.9 53.2 37.6 56.7 28.1 82.3 85.7 75.7 85.0 95.3 79.3 64.6 75.2 58.3
75.7 71.3 77.8 56.0 74.7 46.9 59.8 54.7 62.3 47.2 60.0 40.3 82.0 84.1 77.8 87.6 94.5 84.1 68.0 73.2 64.9
74.8 76.8 73.8 59.9 77.8 51.2 60.0 57.6 61.2 46.4 58.4 40.7 82.7 85.5 77.3 88.4 94.1 85.5 68.7 74.9 64.9
Improv. over baseline +7.1 +0.8 +9.8 +4.8 +3.5 +5.4 +5.7 +0.2 +8.4 +9.9 +3.4 +14.2 +0.8 +0.1 +2.1 +3.9 +0.1 +4.7 +5.3 +1.2 +7.4
Table 3: Performance comparison across different datasets with Supervised and Unsupervised ETF Alignment
5.2 Comparison with State of the Art Methods
We evaluate our NC-GCD framework on several fine-grained and generic classification benchmarks to demonstrate its efficacy in GCD. Tab. 1 presents a comparison between our method and the SOTA methods [3, 4, 5, 10, 11, 6, 8, 7, 40], evaluated on both fine-grained and generic datasets with and without access to the ground-truth (GT) number of classes K for clustering. For the case without access to GT K, we estimate the K through clustering, which is then used to align the ETF prototypes. Meanwhile, to comprehensively evaluate the performance of various methods across Fine-Grained Datasets, Generic Classification Datasets, and All datasets, Table 2 presents the average performance.
Fine-Grained Datasets. On fine-grained benchmarks such as CUB-200, Stanford Cars, FGVC Aircraft, and Herbarium19, NC-GCD consistently outperforms previous methods, whether or not the ground-truth K is provided. As shown in Table 2 (a), when the GT K is available, NC-GCD achieves an average all-class accuracy of 60.3%, surpassing SPT by 3.4% and improving novel class accuracy over SPT by 4.8%. Even without GT K, as shown in Table 2 (b), NC-GCD remains robust, surpassing CMS by 3.3% in all-class accuracy. Notably, on CUB-200, NC-GCD outperforms CMS by 5.9% in all-class accuracy and 7.2% in novel class accuracy, demonstrating strong generalization even without the true number of categories. Fine-grained datasets present higher category confusion due to minimal intra-class variation, and our method effectively mitigates this and enhances class separability.
Generic Classification Datasets. NC-GCD achieves strong performance on standard classification datasets such as CIFAR-100 and ImageNet-100, as shown in Table 2. With known K, NC-GCD attains an all-class accuracy of 88.7% on CIFAR-100 and 94.1% on ImageNet-100, outperforming CMS and SPT. Notably, novel class accuracy improves by 3.0% on CIFAR-100 and 2.9% on ImageNet-100. Even when K must be estimated, NC-GCD maintains its competitive edge, achieving 55.5% all-class accuracy, surpassing CMS by 3.3%, demonstrating its adaptability in generic classification GCD scenarios.
Across both fine-grained and generic datasets, NC-GCD consistently achieves strong performance. As shown in Table 2 (a), with GT K, NC-GCD reaches the highest all-class accuracy of 68.7%, surpassing SPT by 3.0% and improving novel class accuracy by 4.2%. Even without GT K, NC-GCD remains robust, surpassing CMS by 3.1% in all-class accuracy and 4.4% in novel class accuracy (Table 2 (b)). By pre-assigning ETF prototypes, NC-GCD effectively improves class separability, mitigating category confusion and ensuring balanced learning across old and novel categories. These results highlight the robustness and versatility of NC-GCD for real-world GCD applications.
5.3 Ablation Study
The Effectiveness of ETF Alignment. To evaluate the impact of Unsupervised ETF Alignment and Supervised ETF Alignment, we conduct ablation studies with four configurations: (1) Baseline (both disabled), (2) Supervised ETF only, (3) Unsupervised ETF only, and (4) Full model (both enabled). Table 3 presents the results, revealing key insights into each component’s contribution.
Unsupervised ETF Alignment: This component yields the most significant improvements, especially in fine-grained datasets. As shown in Table 3, enabling unsupervised ETF alignment alone leads to substantial gains in novel class accuracy. The model effectively learns robust representations, particularly for novel categories, where differentiation between visually similar classes is crucial.
Supervised ETF Alignment: This component stabilizes and preserves old classes’ accuracy. By aligning labeled features with their corresponding ETF prototypes, it ensures the retention of learned knowledge while maintaining high performance on known categories. Though its contribution to novel class accuracy is smaller than that of unsupervised alignment, it plays a crucial role in preventing forgetting and maintaining model stability.
Synergistic Combination: The full model, integrating both alignments, achieves the highest overall accuracy, with an average all-class accuracy of 68.7%, significantly surpassing the baseline and individual variants. This combination enhances novel class discovery while preserving old class performance. Specifically, the average novel class accuracy increases by 7.4%, and the overall accuracy improves by 5.3%. The results highlight the synergy between supervised and unsupervised alignment, demonstrating the NC-GCD framework’s effectiveness in balancing learning across known and novel categories.
Figure 4: Visualization of Comparison with DINOv1, CMS, and our method on CUB-200.
Figure 5: Visualization of our method training process on CUB-200.
SCM CUB Herbarium 19 ImageNet100
ALL Old Novel ALL Old Novel ALL Old Novel
70.3 71.0 70.0 42.6 54.3 36.3 84.3 88.9 82.0
75.7 71.3 77.8 47.2 60.0 40.3 87.6 94.5 84.1
Improv. +5.3 +0.4 +7.8 +4.5 +5.6 +3.9 +3.3 +5.6 +2.1
Table 4: Ablation study of our SCM module
The Effectiveness of SCM. To evaluate the impact of the Semantic Consistency Matcher (SCM), we conduct ablation studies on CUB-200, ImageNet100, and Herbarium 19. As shown in Table 4, enabling SCM consistently improves performance on all three datasets. On Herbarium 19, which poses greater challenges due to its large number of categories and high intra-class variance, SCM provides more significant improvements, with gains of 4.5% for all classes and 5.6% for novel classes, demonstrating its effectiveness in complex classification tasks. By mitigating clustering fluctuations, SCM ensures stable pseudo-label assignments. These results highlight its crucial role in stabilizing training, particularly in datasets with large class numbers and high intra-class similarity.
The Influence of the ETF Scope Coefficient $\alpha$. The ETF Scope Coefficient $\alpha$ plays a crucial role in determining the number of samples selected for alignment with the ETF prototype. By controlling the percentage of high-confidence samples, $\alpha$ influences the model’s ability to focus on either old or novel classes. As seen in Table 3, when $\alpha$ is too low, the model may not leverage enough high-confidence samples, leading to suboptimal performance. On the other hand, as $\alpha$ increases, novel class accuracy improves significantly, especially in fine-grained datasets where class distinctions are subtle. However, when $\alpha$ exceeds a certain threshold, the model’s focus on new classes comes at the expense of old class accuracy, highlighting the trade-off between the two. The best balance is achieved at $\alpha=0.8$, where both old and novel class accuracy reach optimal values, demonstrating the importance of selecting an appropriate $\alpha$.
Ablation study of ETF loss coefficient $\beta$ and Parameter analysis will be presented in the Appendix.
5.4 Visualization
Comparison with SOTA Methods. As shown in Fig. 4, we compare the embeddings extracted by DINO-V1 (a), CMS (b), and our method (c) for ten classes, with (d) focusing on a subset of three classes from (c). The embeddings from DINO-V1 and CMS show significant overlap between classes, whereas our method clearly separates the classes into distinct clusters. Notably, classes 17, 74, and 104 collapses almost into a single point in (c), which aligns with the Neural Collapse hypothesis. This demonstrates that our method enhances feature discriminability, improving category separability across both known and novel classes.
Training Process. Figure 5 shows the t-SNE visualization of NC-GCD training on CUB-200. At epoch 0, features are highly entangled with severe category overlap. As training progresses, clusters gradually form, improving class separability. By epoch 100, categories are more distinct, and at epoch 200, the feature space aligns with the ETF structure, reinforcing clear decision boundaries. More results and analysis will be presented in the Appendix.
Figure 6: Ablation study of the $\alpha$ on CUB-200.
6 Conclusion and Future Work
We propose NC-GCD, a framework for Generalized Category Discovery that integrates unsupervised and supervised ETF alignment to enhance class separability. Extensive experiments demonstrate that NC-GCD consistently outperforms SOTA methods in both known and novel class accuracy. Ablation studies validate the complementary effects of the two alignment strategies, while t-SNE visualizations highlight its effectiveness in distinguishing similar classes, aligning with the Neural Collapse hypothesis. A promising future research direction is extending our NC-GCD to incremental GCD settings.